40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 25 04:54:21 2013, Masayuki, Update, SUS, ETM violin mode violin1.pdfviolin2.pdf
    Reply  Sun Oct 27 03:41:06 2013, rana, Update, SUS, ETMX violin mode ETMX-Vio3-Ring.png
Message ID: 9299     Entry time: Sun Oct 27 03:41:06 2013     In reply to: 9290
Author: rana 
Type: Update 
Category: SUS 
Subject: ETMX violin mode 

  I thought it would be enough to notch the fundamental and the first harmonic, but sometime tonight the 2nd harmonic at 1892.88 Hz also got rung up.

I made a "Violin3" stopband filter for it and measured its Q using the ole DTT heterodyne secret handshake. Seems much too high to me - it would be nice if someone else would look at this plot and estimate the Q from it.

Turned the PSL HEPA switch back ON - I think its been off for at least a week. I turned the HEPA's variac to 20 after finishing the alignment on the table.

Attachment 1: ETMX-Vio3-Ring.png  119 kB  | Hide | Hide all
ETMX-Vio3-Ring.png
ELOG V3.1.3-